xx NullVars

NullVars

Y.Z. <yosef@raven.phyast.pitt.edu>

January 09 2003

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11nxangxevxoutsidexeq 10Scope: shared from NULLGRIDx01INT

11

11nxangxghostxpts10Scope: shared from NULLGRIDx01INT

11

11nxangxptsxinsidexeq 10Scope: shared from NULLGRIDx01INT

11

11nxangxstencilxsize 10Scope: shared from NULLGRIDx01INT

11

11nxradialxpts 10Scope: shared from NULLGRIDx01INT

11

7 Interfaces

General Implements: nullvars

Grid Variables

7.0.1 PUBLIC GROUPS

11 x Group Names x 10x Variable Names xx01 Details x11x
11 realcharfuncs 10  x01  compact110
11                  10  wcn  x01 description11Evolved Real-valued Characteristic Fields
11                  10  wcs  x01 dimensions112
11                  10  bcn  x01 distribution11DEFAULT
11                  10  bcs  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x01 group type11ARRAY
11                  10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01 timelevels112
11                  10  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  x01variable type11REAL
11cmplxcharfuncsxbasic10  x01  compact110
11                  10  jcn  x01 description11Evolved complex-valued Characteristic Fields (basic set)
11                  10  jcs  x01 dimensions112
11                  10  dxjcn  x01 distribution11DEFAULT
11                  10  dxjcs  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  eth2jcn  x01 group type11ARRAY
11                  10  eth2jcs  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  ucn  x01 timelevels112
11                  10  ucs  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  qcn  x01variable type11COMPLEX
11 cmplxcharfuncsxaux 10  x01  compact110
11                  10  nucn  x01 description11Evolved complex-valued Characteristic Fields (auxiliary set)
11                  10  nucs  x01 dimensions112
11                  10  cbcn  x01 distribution11DEFAULT
11                  10  cbcs  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  ckcn  x01 group type11ARRAY
11                  10  ckcs  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01 timelevels112
11                  10  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  x01variable type11COMPLEX
11 nullxmask 10 x01 compact110
11                  10 boundaryxmaskn x01 description11boundary and evolution masks
11                  10 boundaryxmasks x01 dimensions112
11                  10 evolutionxmaskn x01 distribution11DEFAULT
11                  10 evolutionxmasks x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x01 group type11ARRAY
11                  10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01 timelevels111
11                  10  x01variable type11INT

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullVars. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE